The standard deviation of the market index portfolio is 20%. Stock A has a beta of 1.75 and a residual standard deviation of 30%.

A. calculate the total variance for an increase of 0.25 in its beta

B. Calculate the total variance for an increase of 5.71% in its residual standard deviation.
Answer #1

A. New Beta =1.75+0.25 =2
Total variance =Beta^2*Standard Deviation of Market^2+Residual Standard Deviation^2 =2^2*20%^2+30%^2 =25%

B. New Residual Standard Deviation =30%+5.71% =35.71%
Total variance =Beta^2*Standard Deviation of Market^2+Residual Standard Deviation^2 =1.75^2*20%^2+35.71%^2 =25.00%